OpenStudy (anonymous):

Algebraic Expressions Review.. leave answers in factored form. (x+2)/(x-1)-2/(x+6)-14/(x^2+5x-6) The answer given is x/(x-1). I don't understand how to get that, though.

OpenStudy (anonymous):

one request please write it with equation forum

OpenStudy (anonymous):

\[\frac{x+2}{x-1}\] - \[\frac{2}{x+6}\] - \[\frac{14}{x^2+5x-6}\]

OpenStudy (anonymous):

k

OpenStudy (anonymous):

i got (x^2+6x)/(x^2 + 5x-6) that can be written as x/(x-1) in simplest form .

OpenStudy (anonymous):

do u want process ?

OpenStudy (anonymous):

But, I don't understand how you get that yes please

OpenStudy (anonymous):

ok , see here let me take first of all the first two terms that are : (x+2)/(x-1) - 2/(x+6) now we have to first of all solve this : LCM of (x-1)(x-6) = (x-1)*(x-6) now u can easily solve this : \[[(x+2)(x+6)-2(-1)]\div[(x-1)(x+6)] \] u would get: \[(x^2 + 6x + 14)\div(x^2+5x-6) \]

OpenStudy (anonymous):

now we have this expression \[(x^2+6x+14)/(x^2+5x-6)\] and the other one : \[(14)/(x^2+5x-6)\] \[(x^2+6x+14)/(x^2+5x-6) - (14)/(x^2+5x-6)\] since denominators are same hence we can write it like this: \[{(x^2+6x+14-14)} / (x^2+5x-6)\] that is \[(x^2+6x)/(x^2+5x-6)\]

OpenStudy (anonymous):

now we have this expression \[(x^2+6x+14)/(x^2+5x-6)\] and the other one : \[(14)/(x^2+5x-6)\] \[(x^2+6x+14)/(x^2+5x-6) - (14)/(x^2+5x-6)\] since denominators are same hence we can write it like this: \[{(x^2+6x+14-14)} / (x^2+5x-6)\] that is \[(x^2+6x)/(x^2+5x-6)\]

OpenStudy (anonymous):

now see: x^2+ 6x= x(x+6) and x^2+5x-6 =(x-1)(x-6) that is \[(x^2+6x)/(x^2+5x-6)\] = \[{x(x+6)}/({x-1)(x+6)}\]

OpenStudy (anonymous):

that is \[x/(x-1)\]

OpenStudy (anonymous):

kensey sorry for slow replies my net was disconnecting but finally i had answered :) got it kensey any problem u can ask me now

OpenStudy (anonymous):

THANK YOU SO MUCH!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

a medal would be gr8